Suppose the following bond quotes for IOU Corporation appear in the financial page of today's newspaper. Assume the bond has semiannual payments, a face value of $2,000 and the current date is April 19, 2018. 

Company (Ticker)CouponMaturityLast PriceLast YieldEST Vol (000s) 
 IOU (OU)5.7Apr 19, 2034108.96??1,827

a. What is the yield to maturity of the bond? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

b. What is the current yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 


а. ҮTM = _______ % 

b. Current yield = _______ %
Answer #2

Using financial calculator BA II Plus - Input details:

#

FV = Future Value / Face Value =

-$2,000.00

PV = Present Value = 108.96%*2000 =

$2,179.20

N = Number of years remaining x frequency = 16 x 2 =

32

PMT = Payment = Coupon / frequency =

-$57.00

CPT > I/Y = Rate per period or YTM per period =

            2.443186

a. YTM annual = Periodic YTM*frequency / 100 = 2.443186 x 2 /100

4.89%

b. Current Yield = |PMT|*frequency / PV = 57 x 2 /$2,179.20 =

5.23%
Answer #1

a.

use financial calculator as below

FV=2000

PV=-2000*108.96%=-2179.20

N=16*2=32 semi annual periods

PMT=2000*5.7%/2=57.00

Click CPT

Click I/Y=2.4432% is semi annual rate

YTM=2.4432%*2=4.89%

b.

current yield=(2000*5.7%)/2179.20=5.23%

the above is answer..